Respected Sir / Mam ,
Please help me to solve this question. I will be very grateful for your help .
My question is
A town has a total population of 50,000 . Out of it 28,000 read the newspaper X and 23,000 read Y while 4000 read both the the papers. The number of persons not reading X and Y both is
As the answer given is : 3000
Please provide me the steps for this question.
Thank you

Answer by chessace(471)

No wonder you have issues with this one, it is poorly stated.
The number of persons NOT reading "X and Y both" is 46,000.
It should read "How many persons read neither (no, nadda, zippo) paper?"
Ans: 3000.
Method: The number of papers read is X + Y = 51,000.
The number of persons reading them is less due to some of those persons being counted twice, so subtract them out.
51000 - 4000 = 47000 persons read at least one of those papers.
You are (or should have been) asked How many NOT readers? so that's all the people - those who read:
50000 - 47000 = 3000.
This is a case of the given answer helping to better word the question.